# Chiral Domain Walls Induced by Radially Magnetized Nanotube Geometry

### v1

@@ IV Spin-orbit torque-driven domain wall motion
The dynamics of the DW are described by the following differential equations within the framework of an analytical 1DM [54, 55, 46, 56, 57]:
$\frac{{\rm d}\phi}{{\rm d}t}$ $=\frac{\gamma}{1+\alpha^{2}}\left(-Q\Gamma_{1}-\alpha\Gamma_{2}\right),$ (9)
$\frac{{\rm d}\left(X/\lambda\right)}{{\rm d}t}$ $=\frac{\gamma}{1+\alpha^{2}}\left(-\alpha\Gamma_{1}+Q\Gamma_{2}\right),$
@@
$\Gamma_{2}$ $=Q\xi_{{\rm STT}}j+\frac{1}{2M_{{\rm s}}\lambda}\frac{\partial w}{\partial\phi},$
where $\gamma$ is the gyromagnetic ratio, $\alpha$ is the Gilbert damping, $j$ is the current density, and $w=w_{{\rm DMI}}\left(\phi\right)+w_{{\rm exc}}\left(\phi\right)+w_{{\rm dem}}\left(\phi\right)$ is the energy density per cross-sectional area of the nanotube. The damping-like torque (DLT) efficiency induced by spin Hall effects is given by $\xi_{{\rm DLT}}=\frac{\hbar\theta_{{\rm SH}}}{2M_{{\rm s}}\left|e\right|t}$, and the STT efficiency is defined as $\xi_{{\rm STT}}=-\frac{\hbar P}{2M_{{\rm s}}\left|e\right|\lambda}$. Here, $e$ is the elementary charge, $\hbar$ is the Dirac constant, $\theta_{{\rm SH}}$ is the spin Hall angle (SHA) , and $P$ is the spin polarization. We assume that the same current density is present in both the HM and FM layers [69]. Based on the findings from the previous section, $\Gamma_{2}$ in Eq. (9) can be expressed as
@@
Figure 7 shows the transient position of the DW in the PMA nanotube under applied currents. The results from 1DM and µM simulations are consistent, demonstrating the validity of the approximation that the DW width is constant in the 1DM. The magnetization angle of the DW remains constant during current flow, and the direction of DW motion depends on the sign of $D$. The DWs driven by SOT are slightly displaced from the perfect Bloch-type DW ($\phi=\pi/2$). In systems with a negative SHA, it has been theoretically shown that the DW moves in the direction of current (electron) flow for the right (left)-handed Néel-type DW [81, 69]. Our computational results are consistent with this prediction. These results confirm that the SOT-driven motion in PMA nanotubes is qualitatively similar to that observed in PMA nanoribbons.
@@
Figure 7: Transient position of DW in the PMA nanotube under applied currents. Images are snapshots of the magnetization structure by µM simulations. The enlarged views show the magnetization structure in the vicinity of the DW, with the green arrows indicating the direction of magnetization of the DW. The following parameters are used; $A=1\,{\rm uerg/cm}$, $K_{\rm u}=8.3\,{\rm Merg/cm}^{3}$, $\alpha=0.2$, $P=0.5$, $\theta_{{\rm SH}}=-0.2$, $R_{1}=25\,{\rm nm}$, and $t=1\,{\rm nm}$.
@@
Figure 8 shows the DW velocity as a function of the current density ($v$-$j$ curve) for the PMA nanotube. To clarify the geometrical effects, we also present the results of PMA nanoribbon. In systems where only SOT is present without STT, a comparison of the $v$-$j$ curves for the nanotube and nanoribbon ( Fig. 8(a) and (b)) reveals a more pronounced influence of the unique DMI-like field in the PMA nanotube. In the nanoribbon, the direction of DW motion reverses with a change in the sign of $D$, but the speeds of the DW motion are the same. In contrast, in the nanotube, the direction of DW motion does not necessarily reverse with a change in the sign of $D$, and the speed of DW motion varies accordingly. It should be noted that geometry-induced positive DMI contribution is always present in our PMA nanotube models.
@@
Figure 8: DW velocity as a function of the current density ($v$-$j$ curve) in the PMA nanotube with $R_{1}=25\,{\rm nm}$ and the PMA nanoribbon with a width of $50\,{\rm nm}$. The anisotropy-like field ($H_{\rm a}$) used for the nanoribbons in the 1D simulation was derived from energy calculations obtained through µM simulations. The pulse widths for the calculations are 2 ns. The following parameters are used; $A=1\,{\rm uerg/cm}$, $K_{\rm u}=8.3\,{\rm Merg/cm}^{3}$, $\alpha=0.2$, $\theta_{{\rm SH}}=-0.2$, and $t=1\,{\rm nm}$.
@@
In systems where both STT and SOT are present, SOT serves as the driving force for DW motion, while STT indirectly affects DW motion by modulating the DW magnetization angle [69]. In the case of $D=+0.1\,{\rm erg/cm^{2}}$, the higher DW speed observed in the nanotube (Fig. 8(c)) compared to the nanoribbon (Fig. 8(d)) in the low current density region. This is attributed to the presence of a specific positive-sign DMI-like field within the PMA nanotube. The decrease in DW velocity with increasing current density in the high current region is the characteristic of systems with a negative SHA and a positive DMI [69]. This behavior reflects the transition of the DW from a right-handed Néel-type DW to a left-handed Néel-type DW, driven by increasing STT with higher currents.
@@
In PMA nanotubes, a geometry-induced positive DMI-like field enables DW motion in the direction of the current flow even when $D\leq 0$ and $\theta_{\rm SH}<0$, which differs from PMA nanoribbons [81]. As shown in Fig. 8(d), the DW in the PMA nanoribbon only moves in the direction of electron flow when $D\leq 0$, whereas in the PMA nanotube, the DW moves in the direction of the current flow in the low current density region even when $D\leq 0$ shown in Fig. 8(c). The behavior observed in the PMA nanotube is consistent with the total DMI field, where $H_{{\rm D}}=H_{{\rm D,dem}}^{{\rm tube}}+H_{{\rm DMI}}>0$.
In PMA nanotubes, even when the material-induced DMI is weak, geometrical effects can stabilize a Néel-type DW structure during current flow, enabling higher speed motion. Figure 9 shows the $v$-$j$ curve for PMA nanotube with $D=0$ for different combinations of nanotube radius and exchange stiffness of the magnetic layer. As the radius decreases, both $H_{{\rm D,dem}}^{{\rm tube}}$ and $\left|H_{{\rm a,exc}}^{{\rm tube}}\right|$ increase (Fig. 5). This leads to the stabilization of a right-handed Néel-type DW (Fig. 6), which in turn increases the speed of DW motion. Additionally, an increase in exchange stiffness results in a larger $\left|H_{{\rm a,exc}}^{{\rm tube}}\right|$, further enhancing the stabilization of the Néel-type DW.
@@
Figure 9: Comparison of $v$-$j$ curves for PMA nanotubes with different radius $(R_{1})$ and exchange stiffness constant $(A)$. The pulse widths for the calculations are 5 ns. The following parameters are used; $K_{\rm u}=8.3\,{\rm Merg/cm}^{3}$, $D=0$, $\alpha=0.2$, $P=0.5$, $\theta_{{\rm SH}}=-0.2$, and $t=1\,{\rm nm}$.
@@
$v=-\frac{\gamma\lambda}{\alpha}\frac{\pi}{2}\xi_{\rm DLT}j\left[1-\frac{\left(\alpha^{-1}\frac{\pi}{2}\xi_{\rm DLT}+\xi_{\rm STT}\right)^{2}}{2\left(\frac{\pi}{2}H_{\rm D}-H_{\rm a}\right)^{2}}j^{2}\right].$ (11)
The second term reflects the deceleration that occurs when the DW deviates from the perfect Néel-type DW ($\phi=0,\pi$) during current flow. As the radius of the nanotube increases or the exchange stiffness increases, the denominator of this second term becomes larger (note that $H_{\rm D}>0$ and $H_{\rm a}<0$), resulting in a reduction of the deceleration effect. Consequently, the speed of the DW increases.
@@ V Conclusions
Our findings demonstrate that the stray field originating from the magnetostatic interactions specific to PMA nanotubes can have a contribution comparable to the material-induced DMI for diameters in the sub-$100\,{\rm nm}$ range. Consequently, the direction and speed of DW motion in PMA nanotubes may differ from those observed in nanoribbons composed of the same material.
@@ Appendix A Micromagnetic simulations
We conducted micromagnetic (µM) simulations of a DW inside a PMA nanotube to validate our analytical formulation. Our µM program utilizes the finite volume method [83], and the computational domain includes both ferromagnetic and non-magnetic regions, as shown in Fig. 10(a) and (b). This method is compatible with unstructured meshes, and we therefore employed uniform unstructured meshes in the azimuthal direction to minimize any extraneous effects arising from the mesh geometry. The height of the nanotube in our µM simulation model is $700\,{\rm nm}$ (within the range of $z=-350$ to $+350\,{\rm nm}$). To minimize the impact of end effects on the DW, we conducted shift simulations with pulse width conditions designed to keep the DW within the range of $z=-200$ to $+200\,{\rm nm}$.
@@ Appendix B Derivation of magnetostatic interaction terms
$\frac{\sigma_{{\rm m}}\left(\bm{x}\right)}{M_{\rm s}}$ =$=\begin{cases}-Q\tanh\frac{z-X}{\lambda}&{\rm at}\,r=R_{1}\\ +Q\tanh\frac{z-X}{\lambda}&{\rm at}\,r=R_{0}\\ \end{cases},$ (12)
@@
$E_{{\rm dem,bb}}=$  $\,2\pi^{2}M_{{\rm s}}^{2}\lambda^{3}\int_{0}^{+\infty}{\rm d}\kappa\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho\,J_{0}\left(\rho\kappa\right)\rho\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho^{\prime}J_{0}\left(\rho^{\prime}\kappa\right)\rho^{\prime}\int_{-\infty}^{+\infty}{\rm d}x\int_{-\infty}^{+\infty}{\rm d}y$
×$\times\left(\frac{Q}{\rho}+\cos\phi\,{\rm sech}x\right)\left(\frac{Q}{\rho^{\prime}}+\cos\phi\,{\rm sech}y\right){\displaystyle\tanh x}\,{\displaystyle\tanh}y\,{\rm e}^{-\left|x-y\right|\kappa},$ (14)
$E_{{\rm dem,bs}}=$  $\,4\pi^{2}QM_{{\rm s}}^{2}\lambda^{3}\int_{0}^{+\infty}{\rm d}\kappa\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho\,J_{0}\left(\rho\kappa\right)\rho\left[J_{0}\left(\rho_{0}\kappa\right)\rho_{0}-J_{0}\left(\rho_{1}\kappa\right)\rho_{1}\right]\int_{-\infty}^{+\infty}{\rm d}x\int_{-\infty}^{+\infty}{\rm d}y$
×$\times\left(\frac{Q}{\rho}+\cos\phi\,{\rm sech}x\right){\displaystyle\tanh}x\,{\displaystyle\tanh}y\,{\rm e}^{-\left|x-y\right|\kappa},$ (15)
where $J_{n}\left(x\right)$ is the $n$-th order Bessel function of the first kind. The term $E_{{\rm dem,ss}}$, which is independent of $\phi$, is not included in the calculation as it does not affect the DW dynamics. Equations (14) and (15) include the terms proportional to $\cos\phi$, and Eq. (14) has the term proportional to $\cos\phi^{2}$. By dividing these by the tube cross-sectional area, the linear and quadratic terms of $\cos\phi$ can be expressed as
$w_{1}=$ $\,Q\frac{4\pi M_{{\rm s}}^{2}\lambda}{\rho_{1}^{2}-\rho_{0}^{2}}\cos\phi\int_{0}^{+\infty}{\rm d}\kappa\left[\rho_{1}J_{1}\left(\rho_{1}\kappa\right)-\rho_{0}J_{1}\left(\rho_{0}\kappa\right)\right]\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho\,J_{1}\left(\rho\kappa\right)\rho\int_{-\infty}^{+\infty}{\rm d}x\,{\displaystyle\tanh}x\int_{-\infty}^{+\infty}{\rm d}y\,\frac{{\tanh}y}{\cosh y}\,{\rm e}^{-\left|x-y\right|\kappa},$ (16)
$w_{2}=$ $\,\frac{2\pi M_{{\rm s}}^{2}\lambda}{\rho_{1}^{2}-\rho_{0}^{2}}\cos^{2}\phi\int_{0}^{+\infty}{\rm d}\kappa\left[\frac{\rho_{1}J_{1}\left(\rho_{1}\kappa\right)-\rho_{0}J_{1}\left(\rho_{0}\kappa\right)}{\kappa}\right]^{2}\int_{-\infty}^{+\infty}{\rm d}x\,\frac{{\tanh}x}{\cosh x}\int_{-\infty}^{+\infty}{\rm d}y\,\frac{{\tanh}y}{\cosh y}\,{\rm e}^{-\left|x-y\right|\kappa}.$ (17)
By comparing each result with Eqs. (3) and (4), we obtain Eqs. (6) and (7).

### v2

@@ IV Spin-orbit torque driven domain wall motion
The dynamics of DWs are described by the following differential equations within the framework of an analytical 1DM [61, 62, 52, 63, 64]:
$\frac{{\rm d}\phi}{{\rm d}t}$ $=\frac{\gamma_{0}}{1+\alpha^{2}}\left(-Q\Gamma_{1}-\alpha\Gamma_{2}\right),$ (9)
$\frac{{\rm d}\left(X/\lambda\right)}{{\rm d}t}$ $=\frac{\gamma_{0}}{1+\alpha^{2}}\left(-\alpha\Gamma_{1}+Q\Gamma_{2}\right),$
@@
$\Gamma_{2}$ $=Q\xi_{{\rm STT}}j+\frac{1}{2\mu_{0}M_{{\rm s}}\lambda}\frac{\partial w}{\partial\phi},$
where $\gamma_{0}$ is the gyromagnetic ratio, $\alpha$ is the Gilbert damping, $j$ is the current density, and $w=w_{{\rm DMI}}\left(\phi\right)+w_{{\rm exc}}\left(\phi\right)+w_{{\rm dem}}\left(\phi\right)$ is the energy density per cross-sectional area of the nanotube. The damping-like torque (DLT) efficiency induced by spin Hall effects is given by $\xi_{{\rm DLT}}=\frac{\hbar\theta_{{\rm SH}}}{2\mu_{0}M_{{\rm s}}\left|e\right|t}$, and the STT efficiency is defined as $\xi_{{\rm STT}}=-\frac{\hbar P}{2\mu_{0}M_{{\rm s}}\left|e\right|\lambda}$. Here, $e$ is the elementary charge, $\hbar$ is the Dirac constant, $\theta_{{\rm SH}}$ is the spin Hall angle (SHA) , and $P$ is the spin polarization. We assume that the same current density is present in both the HM and FM layers [76]. Based on the findings from the previous section, $\Gamma_{2}$ in Eq. (9) can be expressed as
@@
Figure 7 shows the transient position of the DW in the PMA nanotube under applied currents. The results from 1DM and µM simulations are consistent, demonstrating the validity of the approximation that the DW width is constant in the 1DM. The magnetization angle of the DW remains constant during current flow, and the direction of DW motion depends on the sign of $D$. The DWs driven by SOT are slightly displaced from the perfect Bloch-type DW ($\phi=\pi/2$). In systems with a negative SHA, it has been theoretically shown that the DW moves in the direction of current (electron) flow for the right (left)-handed Néel-type DW [90, 76]. Our computational results are consistent with these predictions. These results confirm that the SOT-driven motion in PMA nanotubes is qualitatively similar to that observed in PMA flat nanoribbons.
@@
Figure 7: Transient position of DW in the PMA nanotube under applied currents. Images are snapshots of the magnetization structure by µM simulations. The enlarged views show the magnetization structure in the vicinity of the DW, with the green arrows indicating the direction of magnetization of the DW. The following parameters are used; $A=10^{-11}\,{\rm J/m}$, $K_{\rm u}=0.83\times 10^{6}\,{\rm J/m^{3}}$, $\alpha=0.2$, $P=0.5$, $\theta_{{\rm SH}}=-0.2$, $R_{1}=25\,{\rm nm}$, and $t=1\,{\rm nm}$.
@@
The impact of geometrical effects on SOT-driven DW motion in PMA nanotubes is clarified by comparing it to the DW velocity as a function of current density ($v$-$j$ curve) in PMA flat nanoribbons, as shown in Fig. 8. In systems where only SOT is present and STT is absent, the direction of DW motion in the flat nanoribbon reverses with a change in the sign of $D$, but the speeds of DW motion are the same (Fig. 8(a)). In contrast, in the nanotube, the direction of DW motion does not necessarily reverse with a change in the sign of $D$, and the speeds of DW motion vary accordingly (Fig. 8(b)). It should be noted that a geometry-induced positive DMI contribution is always present in our PMA nanotube models.
@@
Figure 8: DW velocity as a function of current density ($v$-$j$ curve) in the PMA flat nanoribbon with a width of $50\,{\rm nm}$ and the PMA nanotube with $R_{1}=25\,{\rm nm}$. The anisotropy-like field ($H_{\rm a}$) used for the flat nanoribbons in the 1D simulation was derived from energy calculations obtained through µM simulations. The pulse widths for the calculations are 2 ns. The following parameters are used; $A=10^{-11}\,{\rm J/m}$, $K_{\rm u}=0.83\times 10^{6}\,{\rm J/m^{3}}$, $\alpha=0.2$, $\theta_{{\rm SH}}=-0.2$, and $t=1\,{\rm nm}$.
@@
In systems where both STT and SOT are present, SOT serves as the driving force for DW motion, while STT indirectly affects DW motion by modulating the DW magnetization angle [76]. In the case of $D=+0.1\,{\rm mJ/m^{2}}$, the higher DW speed observed in the nanotube (Fig. 8(d)) compared to the flat nanoribbon (Fig. 8(c)) in the low current density region. This is attributed to the presence of a specific positive-sign DMI-like field within the PMA nanotube. The decrease in DW velocity with increasing current density in the high current region is the characteristic of systems with a negative SHA and a positive DMI [76]. This behavior reflects the transition of the DW from a right-handed Néel-type DW to a left-handed Néel-type DW, driven by increasing STT with higher currents.
@@
In PMA nanotubes, a geometry-induced positive DMI-like field enables DW motion in the direction of the current flow even when $D\leq 0$ and $\theta_{\rm SH}<0$, which differs from PMA flat nanoribbons [90]. As shown in Fig. 8(c), the DW in the PMA flat nanoribbon only moves in the direction of electron flow when $D\leq 0$, whereas in the PMA nanotube, the DW moves in the direction of the current flow in the low current density region even when $D\leq 0$ shown in Fig. 8(d). The behavior observed in the PMA nanotube is consistent with the total DMI field, where $H_{{\rm D}}=H_{{\rm D,dem}}^{{\rm tube}}+H_{{\rm DMI}}>0$.
Geometrical effects in PMA nanotubes can stabilize a Néel-type DW structure during current flow, even when the material-induced DMI is weak, enabling highly efficient SOT-driven DW motion. Figure 9 shows that structures with smaller radii and materials with larger exchange stiffness exhibit higher SOT efficiency. The reduction of radius and the increase in exchange stiffness promote the stabilization of Néel-type DWs, as they increase the exchange energy associated with the Bloch-type DWs (which corresponds to the decrease in $H_{{\rm a,exc}}^{{\rm tube}}$ with reduced $R_{1}$ and increased $A$, as shown in Fig. 5(a)). Additionally, the reduction in radius enhances the stray field from the magnetic domains acting on the DW (which corresponds to the increase in $H_{{\rm D,dem}}^{{\rm tube}}$ with reduced $R_{1}$, as shown in Fig. 5(c)), further stabilizing right-handed Néel-type DWs. These combined effects suggest that tailoring geometric parameters can effectively induce and stabilize Néel-type DWs, thereby achieving higher-speed SOT-driven DW motion without relying solely on the material-induced DMI. Furthermore, to achieve the same speed of DW motion, a lower current density is required in PMA nanotubes compared to PMA flat nanoribbons.
@@
Figure 9: Comparison of $v$-$j$ curves for PMA nanotubes with different radius $(R_{1})$ and exchange stiffness constant $(A)$. The pulse widths for the calculations are 5 ns. The following parameters are used; $K_{\rm u}=0.83\times 10^{6}\,{\rm J/m^{3}}$, $D=0$, $\alpha=0.2$, $P=0.5$, $\theta_{{\rm SH}}=-0.2$, and $t=1\,{\rm nm}$.
@@
$v=-\frac{\gamma_{0}\lambda}{\alpha}\frac{\pi}{2}\xi_{\rm DLT}j\left[1-\frac{\left(\alpha^{-1}\frac{\pi}{2}\xi_{\rm DLT}+\xi_{\rm STT}\right)^{2}}{2\left(\frac{\pi}{2}H_{\rm D}-H_{\rm a}\right)^{2}}j^{2}\right].$ (11)
The second term reflects the deceleration that occurs when the DW deviates from the perfect Néel-type DW ($\phi=0,\pi$) during current flow. As the radius of the nanotube decreases or the exchange stiffness increases, $H_{{\rm a}}=H_{{\rm a,exc}}^{{\rm tube}}+H_{{\rm a,dem}}^{{\rm tube}}$ increases in the negative direction, while $H_{{\rm D}}=H_{{\rm D,dem}}^{{\rm tube}}+H_{{\rm DMI}}$ increases in the positive direction (see Fig. 5), resulting in a reduction of the deceleration effect. In flat nanoribbons, which represent the limiting case where the radius is infinitely large, the deceleration effect is more pronounced compared to nanotubes, leading to slower DW motion.
@@ V Conclusions
Our findings demonstrate that the stray field originating from the magnetostatic interactions specific to PMA nanotubes can have a contribution comparable to the material-induced DMI for diameters in the sub-$100\,{\rm nm}$ range. Consequently, the direction and speed of DW motion in PMA nanotubes may differ from those observed in flat nanoribbons composed of the same material.
@@ Appendix A Micromagnetic simulations
We conducted micromagnetic (µM) simulations [63] of a DW inside a PMA nanotube to validate our analytical formulation. Our µM program utilizes the finite volume method [92], and the computational domain includes both ferromagnetic and non-magnetic regions, as shown in Fig. 10(a) and (b). This method is compatible with unstructured meshes, and we therefore employed uniform unstructured meshes in the azimuthal direction to minimize any extraneous effects arising from the mesh geometry. The height of the nanotube in our µM simulation model is $700\,{\rm nm}$ (within the range of z=350z=-350 to $350\,{\rm nm}$). To minimize the impact of end effects on the DW, we conducted shift simulations with pulse width conditions designed to keep the DW within the range of z=200z=-200 to $200\,{\rm nm}$.
@@ Appendix B Derivation of magnetostatic interaction terms
$\frac{\sigma_{{\rm m}}\left(\bm{x}\right)}{M_{\rm s}}$ $=\begin{cases}-Q\tanh\frac{z-X}{\lambda}&{\rm at}\,r=R_{1}\\ +Q\tanh\frac{z-X}{\lambda}&{\rm at}\,r=R_{0}\\ \end{cases},$ (12)
@@
$E_{{\rm dem,bb}}=$ $\,\frac{\pi}{2}\mu_{0}M_{{\rm s}}^{2}\lambda^{3}\int_{0}^{+\infty}{\rm d}\kappa\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho\,J_{0}\left(\rho\kappa\right)\rho\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho^{\prime}J_{0}\left(\rho^{\prime}\kappa\right)\rho^{\prime}\int_{-\infty}^{+\infty}{\rm d}x\int_{-\infty}^{+\infty}{\rm d}y$
$\times\left(\frac{Q}{\rho}+\cos\phi\,{\rm sech}x\right)\left(\frac{Q}{\rho^{\prime}}+\cos\phi\,{\rm sech}y\right){\displaystyle\tanh x}\,{\displaystyle\tanh}y\,{\rm e}^{-\left|x-y\right|\kappa},$ (14)
$E_{{\rm dem,bs}}=$ $\,\pi Q\mu_{0}M_{{\rm s}}^{2}\lambda^{3}\int_{0}^{+\infty}{\rm d}\kappa\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho\,J_{0}\left(\rho\kappa\right)\rho\left[J_{0}\left(\rho_{0}\kappa\right)\rho_{0}-J_{0}\left(\rho_{1}\kappa\right)\rho_{1}\right]\int_{-\infty}^{+\infty}{\rm d}x\int_{-\infty}^{+\infty}{\rm d}y$
$\times\left(\frac{Q}{\rho}+\cos\phi\,{\rm sech}x\right){\displaystyle\tanh}x\,{\displaystyle\tanh}y\,{\rm e}^{-\left|x-y\right|\kappa},$ (15)
where $J_{n}\left(x\right)$ is the $n$-th order Bessel function of the first kind. The term $E_{{\rm dem,ss}}$, which is independent of $\phi$, is not included in the calculation as it does not affect the DW dynamics. Equations (14) and (15) include the terms proportional to $\cos\phi$, and Eq. (14) has the term proportional to $\cos^{2}\phi$. By dividing these by the tube cross-sectional area, the linear and quadratic terms of $\cos\phi$ can be expressed as
$w_{1}=$ $\,Q\frac{\mu_{0}M_{{\rm s}}^{2}\lambda}{\rho_{1}^{2}-\rho_{0}^{2}}\cos\phi\int_{0}^{+\infty}{\rm d}\kappa\left[\rho_{1}J_{1}\left(\rho_{1}\kappa\right)-\rho_{0}J_{1}\left(\rho_{0}\kappa\right)\right]\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho\,J_{1}\left(\rho\kappa\right)\rho\int_{-\infty}^{+\infty}{\rm d}x\,{\displaystyle\tanh}x\int_{-\infty}^{+\infty}{\rm d}y\,\frac{{\tanh}y}{\cosh y}\,{\rm e}^{-\left|x-y\right|\kappa},$ (16)
$w_{2}=$ $\,\frac{1}{2}\frac{\mu_{0}M_{{\rm s}}^{2}\lambda}{\rho_{1}^{2}-\rho_{0}^{2}}\cos^{2}\phi\int_{0}^{+\infty}{\rm d}\kappa\left[\frac{\rho_{1}J_{1}\left(\rho_{1}\kappa\right)-\rho_{0}J_{1}\left(\rho_{0}\kappa\right)}{\kappa}\right]^{2}\int_{-\infty}^{+\infty}{\rm d}x\,\frac{{\tanh}x}{\cosh x}\int_{-\infty}^{+\infty}{\rm d}y\,\frac{{\tanh}y}{\cosh y}\,{\rm e}^{-\left|x-y\right|\kappa}.$ (17)
Using the identity $\int_{-\infty}^{+\infty}{\rm d}x\,f(x)\int_{-\infty}^{+\infty}{\rm d}y\,\frac{{\tanh}y}{{\cosh}y}{\rm e}^{-\left|x-y\right|\kappa}=2\kappa\int_{0}^{+\infty}{\rm d}y\,{\rm e}^{-y\kappa}\int_{-\infty}^{+\infty}{\rm d}x\,f(x)/\cosh(y-x)$, which holds for the odd function $f(x)$, and integrating both Eq. (16) and Eq. (17) with respect to $x$, we can compare the results to Eqs. (3) and (4), respectively, to obtain Eqs. (6) and (7) .